# When Do Tools and Planning Help LLMs Think? A Cost- and Latency-Aware Benchmark

### v1

@@ IV-C2 ChangeMyView results
Figure 6: ChangeMyView - GPT-4o
Figure 7: ChangeMyView - GPT-4o-mini
@@
Figure 8: ChangeMyView - Best Overall
TABLE III: ChangeMyView: Approach and Model Performance Comparison Across Accuracy, Speed, and Configuration
@@ V Discussion
In EventQA, which involves complex data interpretation and analysis, GPT-4o appears to be better adapted at handling complex multi-stage thought processes as well as more complicated tools. In the DBpedia approach, the DBpedia queries required an element of graph schema discovery, utilizing tools provided to deduce and learn a graph ontology. GPT-4o achieved overall superior results by being able to utilize the tools and mult-stage approaches more efficiently.
@@ VI Conclusion
This paper introduced an LLM Thinking Benchmark for studying how planning and tool use affect real-world performance under practical cost and latency constraints. We evaluated two representative settings: (i) event-centric question answering over graph-structured knowledge (Event-QA), and (ii) persuasive response generation in open-domain discussions (ChangeMyView). Across both settings, we compared a fast one-shot baseline against a plan–execute–replan agent implemented with LangGraph and task-specific tools.
@@
In contrast, for ChangeMyView the simplest approach was consistently strongest: the NoPlanning baseline (especially with GPT-4o-mini) achieved the best overall accuracy with minimal latency (Table III). The PlanningSearch configuration often increased runtime substantially without improving final accuracy, indicating that additional tool calls and deliberation can be counterproductive on tasks where a model’s internal priors already align well with the expected response style and where retrieval adds noise or distracts from argument quality.

### v2

@@ IV-C2 CMV Results
Figure 6: CMV - GPT-4o
Figure 7: CMV - GPT-4o-mini
@@
Figure 8: CMV - Best Overall - NoPlanning (GPT-4o-mini, 75% accuracy at  6s latency).
TABLE III: CMV: Aggregate view of the best-performing configuration per model relative to the baseline, comparing accuracy, speed, and configuration.
@@ V Discussion
In Event-QA, which involves complex data interpretation and analysis, GPT-4o appears to be better adapted at handling complex multi-stage thought processes as well as more complicated tools. In the DBpedia approach, the DBpedia queries required an element of graph schema discovery, utilizing tools provided to deduce and learn a graph ontology. GPT-4o achieved overall superior results by being able to utilize the tools and multi-stage approaches more efficiently.
@@ VI Conclusion
This paper introduced an LLM Thinking Benchmark for studying how planning and tool use affect real-world performance under practical cost and latency constraints. We evaluated two representative settings: (i) event-centric question answering over graph-structured knowledge (Event-QA), and (ii) persuasive response generation in open-domain discussions (CMV). Across both settings, we compared a fast one-shot baseline against a plan–execute–replan agent implemented with LangGraph and task-specific tools.
@@
In contrast, for CMV the simplest approach was consistently strongest: the NoPlanning baseline (especially with GPT-4o-mini) achieved the best overall accuracy with minimal latency (Table III). The PlanningSearch configuration often increased runtime substantially without improving final accuracy, indicating that additional tool calls and deliberation can be counterproductive on tasks where a model’s internal priors already align well with the expected response style and where retrieval adds noise or distracts from argument quality.